OpenStudy (anonymous):

how do you solve x+3/6 = 3/8 + x-5/4 (the slashes mean over)

OpenStudy (anonymous):

is that (x + 3)/6 or x + (3/6) ? similar for x-5/4?

OpenStudy (anonymous):

(x+3) over 6 = 3/8 + (x-5) over 4

OpenStudy (anonymous):

let's multiply everything by 24 to make the math easy: 4x + 12 = 9 + 6x - 30 4x + 3 = 6x - 30 3 = 2x - 30 2x = 33 x = 33/2

OpenStudy (anonymous):

thank you !